OpenStudy (anonymous):

Find the derivative of this function g(x) = x^2-2x + 5/x This is what I got f(x) = (X^2 - 2X +5) f' (x) = (2x -2 +5) g(x) = (x) g'(x) = 1 g'= (2x -2 +5) (x) -1 (x^2- 2x +5)/ (x+1) Get you help to tell me if I am in the right direction.....

OpenStudy (anonymous):

Rewrite it as g(x)= x^2 -2x + 5x^-1

OpenStudy (anonymous):

\[\frac{ x^2-2x + 5 }{ x }\] Need to find the derivative of this problem

OpenStudy (anonymous):

Simplify it or do the quotient rule. I would do (x^2/ x) - (2x/x) + 5/x which ends as x-2+5x^-1

OpenStudy (anonymous):

g(x) = x-2+5x^-1 g'(x) = 1-5x^-2 g'(x) = 1-(5/x^-2)